OpenStudy (anonymous):

Eliminating the Arbitrary Constants in DE:

OpenStudy (anonymous):

\[y=c _{1}+c _{2}e ^{-4x}\]

OpenStudy (anonymous):

so i eliminated the arbitrary constant and i just want to check if my answer is correct. which is \[y \prime \prime=-4\]

OpenStudy (unklerhaukus):

what is y'

OpenStudy (anonymous):

\[y \prime=-4c _{2}e ^{-4x}\]

OpenStudy (jhannybean):

\[\large y'= 0-[0\cdot e^{-4x} -4e^{-4}\cdot c_{2}] \]\[\large y'= -4c_{2}e^{-4x}\]

OpenStudy (unklerhaukus):

good

OpenStudy (jhannybean):

I think you can then take \(-4c_{2}\) as f(x) and \(e^{-4x}\) as g(x) and apply the product rule again.

OpenStudy (jhannybean):

...and i forgot the "x" in my first line of my last post. Supposed to be \(\large -4e^{-4x}\)

OpenStudy (anonymous):

this is what i did

OpenStudy (anonymous):

is that right? :/

OpenStudy (jhannybean):

Hmm...so \[\large y'' = [0\cdot e^{-4x} +[-4e^{-4x}\cdot (-4c_{2})]]\]\[\large y''= 0+(-4)(-4)c_{2}e^{-4x}\]\[\large y'' = 0+16c_{2}e^{-4x}\]

OpenStudy (anonymous):

@kenji07 I agree with @Jhannybean work above.

OpenStudy (jhannybean):

Maybe i'm not understanding your steps in eliminating the constants. i just used the product rule but it's not an efficient way of getting rid of the constants. Instead of writing out what you did,can you explain to us step by step what you did to solve your problem?

OpenStudy (anonymous):

i did that first then changed my mind. how can i eliminate Csub2?

OpenStudy (jhannybean):

That's why i'm asking you to explain what you did to get rid of \(c_{1}\) and \(c_{2}\)

OpenStudy (anonymous):

i attached it :)

OpenStudy (anonymous):

i just manipulated the equations :\

OpenStudy (jhannybean):

I see tht you've tken thederivative to get rid of the first constant..... what did you do in the third step? \[\large y-c_{1}=c_{2}e^{-4x}\]?

OpenStudy (anonymous):

manipulate the original equation to get c2(e^−4x)

OpenStudy (anonymous):

c2(e^−4x)=y-c1. this is my third equation. i sustituted this in the first derivative

OpenStudy (jhannybean):

I have a feeling that's not a valid manipulation :\ Basically you're saying that you have \[\large y= c_{1} + c_{2}e^{-4x}\] then \[\large y-c_{1}=c_{2}e^{-4x}\]\[\large \frac{dy}{dx}(y-c_{1}=c_{2}e^{-4x})\]\[\large y' = -4c_{2}e^{-4x}\] So far....

OpenStudy (jhannybean):

Oh scratch out my first comment... bleh.

OpenStudy (anonymous):

what could i do to eliminate c2 after getting te second derivative :/

OpenStudy (jhannybean):

Because now i'm not understanding your 4rth step. \[\large \text{sub 3 in 2} :y'=4(y-c_{1})\]

OpenStudy (anonymous):

you can't get rid of all constants whatsoever

OpenStudy (jhannybean):

^ that's what i was thinking.

OpenStudy (anonymous):

there is a negative before 4

OpenStudy (jhannybean):

Sorry about that.

OpenStudy (jhannybean):

You're stating that the function, \(y\) is dependent on the rate of change \(y'\) which is totally false.

OpenStudy (jhannybean):

the rate of change change, \(y'\), is dependent on the function \(y\) and therefore your fourth step does not make sense.

OpenStudy (anonymous):

ok. thanks. how about this one?

OpenStudy (anonymous):

@kenji07 That is correct.

OpenStudy (jhannybean):

That is correct. you found all your derivatives first, and found that the orignial equation equalled the same value as your original function,therefore \(y'' =y\)

OpenStudy (anonymous):

im going to bed gn guys

OpenStudy (anonymous):

ok. thanks guys :)

OpenStudy (jhannybean):

Ask yourself, "What does \(y''\) mean in relation to the function \(y \)?"

OpenStudy (anonymous):

hey. uhm. how about this one? did i get it right?

OpenStudy (anonymous):

@Jhannybean ?

OpenStudy (jhannybean):

No that doesn't seem right.how are you factoring out a 2+3 when those arent the LCM of both terms?

OpenStudy (jhannybean):

It seems like you're trying to combine y' and y'' again and that cannot work,when you find the derivative of something you cant constantly switch between derivatives.

OpenStudy (jhannybean):

What exactly are you trying to accomplish here?

OpenStudy (anonymous):

oh. okay. can you teach me how to solve this one?

OpenStudy (anonymous):

still eliminating arbitrary constant

OpenStudy (jhannybean):

\[\large y=c_{1}e^{2x}+c_{2}e^{3x}\]Product rule on both. \[\large y'= [0\cdot e^{2x}+2c_{1}e^{2x}] + [0\cdot e^{3x} + 3c_{2}e^{3x}] \] simplify and use the product rule to get rid of as many terms as you can.

OpenStudy (anonymous):

\[y \prime= 2c _{1}e ^{2x}+3c _{2}e ^{3x}\]

OpenStudy (jhannybean):

I am not going to solve the remaining derivatives for you. Just follow the steps i've taken in the other problems ive helped you solve.

OpenStudy (whpalmer4):

Also, by the logic shown on the paper, you had \[y = c_1 e^{2x} + c_2e^{3x}\]and then you waved your hands a bit and got \[y' = 5y\]but if that is true, then \[y' = 5y = 5*c_1e^{2x} + 5*c_2e^{3x}\]but differentiating \(y\) gives you \[y' = c_1e^{2x} + c_2e^{3x}\]so to match up terms, \[5c_1 = c_1\]and\[5c_2 = c_2\]which is only true if they are 0, and that's not a very interesting diffyq...

OpenStudy (anonymous):

\[y \prime \prime=4c _{1}e ^{2x}+9c _{2}\]

OpenStudy (whpalmer4):

uh, sorry, I forgot the 2 and 3 in front after differentiating, but you still end up with the same bottom line, c_1 = c_2 = 0

OpenStudy (anonymous):

forgot the e^3x

OpenStudy (anonymous):

please help. i really cant get it :(

OpenStudy (anonymous):

\[y \prime \prime - y = 0\] \[y = c _{1}e ^{2x} + c _{2}e^{3x}\] \[y \prime = 2c _{1}e ^{2x} + 3c _{2}e^{3x}\] \[y \prime \prime = 4c _{1}e ^{2x} + 9c _{2}e^{3x}\] \[y \prime \prime - y = 4c _{1}e ^{2x} + 9c _{2}e^{3x} - (c _{1}e ^{2x} + c _{2}e^{3x}) = 3c _{1}e ^{2x} + 8c _{2}e^{3x}\]

OpenStudy (anonymous):

But, since y'' - y = 0, 3c1e^2x + 8c2e^3x = 0 also!

OpenStudy (anonymous):

Do you have an initial value for this problem? What exactly are we trying to do?

OpenStudy (anonymous):

Oh, I see, we can substitute 3y for 3c1e^2x + 3c2e^3x

OpenStudy (anonymous):

3c1e^2x + 8c2e^3x = 0 3y + 5c2e^3x = 0 y = (-5/3)c2e^3x

OpenStudy (anonymous):

\[y = -\frac{ 5 }{ 3 }c _{2}e^{3x}\]

OpenStudy (anonymous):

Still, without an initial value I'm not sure what I am doing. I guess we could go back to 3c1e^2x + 8c2e^3x = 0 and solve for x in terms of c1 and c2.

OpenStudy (anonymous):

3c1e^2x + 8c2e^3x = 0 3c1e^2x = -8c2e^3x 3c1/(-8c2) = e^x ln(-3c1/8c2) = x x = ln(-3c1/8c2)

OpenStudy (anonymous):

Kenji, what was the original problem's instructions?

OpenStudy (anonymous):

@cebroski bro \[y \prime \prime-y=0\] is an answer to another problem. its not related to y=c1e2x+c2e3x.

OpenStudy (anonymous):

im eliminating the arbitrary constant in y=c1e2x+c2e3x. that the original problem's instruction.

OpenStudy (anonymous):

y=c1e2x+c2e3x y'=2c1e2x+3c2e3x y''=4c1e2x+9c2e3x y'=2y+c2e3x c2=(y'-2y)/e3x y''=4c1e2x+9((y'-2y)/e3x)e3x y''=4c1e2x+9(y'-2y) c1=(y''-9(y'-2y))/4e2x y=((y''-9(y'-2y))/4e2x)e2x+((y'-2y)/e3x)e3x y=((y''-9(y'-2y))/4)+y'-2y 4(3y-y')=y''-9(y'-2y) 12y-4y'=y''-9y'-18y 30y+5y'-y''=0 y''-5y'-30y=0